HUBBLE SPACE TELESCOPE DAILY REPORT #2226 PERIOD COVERED: 0000Z (UTC) 10/08/98 - 0000Z (UTC) 10/09/98 Daily Status Report as of 282/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around the main sequence star Alpha-Lyra (a.k.a., Vega). The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.4 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star HD4628. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.5 Completed STIS/CCD/MA2/WF/PC-2/NIC/1/2/3 8076 (HDF-South Observations) The WF/PC-2, The Space Telescope Imaging Spectrograph (MA2 and CCD) and the Near Infrared Camera (Cameras One, Two and Three) all were used to make deep observations of the Hubble Deep Fields- South (HDFS-22333). This program is designed to provide deep observations of a high-galactic latitude field in the southern continuous viewing zone. The STIS field is centered on a bright QSO at z=2.25. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Operations Notes: The Hubble Deep Field - South (HDF-S) campaign continues. There were no problems with any HDF-S observations this operational day. Using ROP SR-1, the SSR EDAC error counter was cleared three times. 3.0 SIGNIFICANT FORTHCOMING EVENTS: This evening, at approximately 6:00 p.m., the flight operations team is expected to transfer operations to the SAMS backup facility. When this occurs, all PRS/legacy systems and other related systems in the HST operations area will be shut down in preparation for scheduled PDU power work. Following the PDU work, operational systems will be powered up with HST operations personnel expected to return to the MOR during the day on Sunday, prior to the mirror move scheduled for approximately 10:00 p.m. Continuation of normal science observations and calibrations. /DMH